OpenStudy (lukecrayonz):

@amistre64 You're being called in for this question.. it's just like the other one -_-

OpenStudy (bahrom7893):

amistre's busy i'll help

OpenStudy (lukecrayonz):

Find an equation of the ellipse with its center at the origin. Foci: (positive & negative 2,0) Major axis length of 8

OpenStudy (lukecrayonz):

2a=8, a=4.

OpenStudy (amistre64):

half the distance between foci is measured as "c" traditionally

OpenStudy (bahrom7893):

nevermind he's not busy..

OpenStudy (amistre64):

and half the major axis as "a"

OpenStudy (amistre64):

\[(x/a)^2+(y/b)^2=1\] and a and b can move under wither numerator depending on how the ecclps is spoes to stretch

OpenStudy (lukecrayonz):

So is that 4^2-b^2=2^2?

OpenStudy (lukecrayonz):

16-b^2=4, b^2=12

OpenStudy (amistre64):